Tiny and Tim each own half of the 100 outstanding shares of Flower Corporation. This year, Flower reported taxable income of $6,000. In addition, Flower received $20,000 of life insurance proceeds due to the death of an employee (Flower paid $500 in life insurance premiums this year). Flower had $5,000 of accumulated E&P at the beginning of the year. (Leave no answer blank. Enter zero if applicable.)

b. Flower distributed $6,000 on February 15 and $30,000 on August 1. What total amount of dividends will Tiny and Tim report? (Do not round intermediate calculations.)

The current E&P is $24,240 !
Answer #1

Current E&P = $24,240

Calculation of total amount of dividends will Tiny and Tim report:

Particulars Tiny Tim
Amount distributed 6,000 30,000
Proportion of total distribution 0.16667 0.83333
Current E & P 4,040 20,200
Add: Accumulated E& P 2,000 3,000
Total dividend 6,040 23,200

Calculation of Proportion of Distribution:-

Tiny = 6000 / 6000 + 30000 = 0.16667

Tim = 30000 / 6000 + 30000 = 0.83333

Calculation of Current E & P:-

Tiny = Total current E&P * Tim's proportion = 24240 * 0.16667 = 4040

Tim = Total current E&P * Tiny's proportion = 24240 * 0.83333 = 20200.

OR

Current E&P =Taxable income + Tax exempt -Federal income tax-expenses

=$ 6000 + $ 20,000-$ 1,500-$ 500=$ 24,000

Calculation of total amount of dividends will Tiny and Tim report:

Particulars Tiny Tim
Amount distributed 6,000 30,000
Proportion of total distribution 0.16667 0.83333
Current E & P 4,000 20,000
Add: Accumulated E& P 2,000 3,000
Total dividend 6,000 23,000

Calculation of Proportion of Distribution:-

Tiny = 6000 / 6000 + 30000 = 0.16667

Tim = 30000 / 6000 + 30000 = 0.83333

Calculation of Current E & P:-

Tiny = Total current E&P * Tim's proportion = 24000 * 0.16667 = 4000

Tim = Total current E&P * Tiny's proportion = 24000 * 0.83333 = 20000.

Please give a Thumbs up ?. Thanks!!